OpenStudy (anonymous):

Find x value for the point that splits segment CD in half if point C is located at (−2, 4) and point D is located at (3, 7).

OpenStudy (anonymous):

hey sweetsunray

OpenStudy (anonymous):

the x distance = 3 - (-2) = ? the y distance = 7-4 = ? pythagoras says that of a square triangle the hypothenuse (the long side) = \[\sqrt{x ^{2}+y ^{2}}\] plug the distances you calculated in the squareroot for x and y, calculate their squares, and it up and then take the root.

OpenStudy (anonymous):

ok im just want you guys to know that this is a pre test im doing for flvs for geometry for the second semester. your gonna be seeing me for a couple more questions